# The separating Noether number of small groups

### v2

@@ 8.4. The group $\mathrm{M}_{27}$
is the non-abelian group of order $27$ with an index $3$ cyclic subgroup. Assume that $K$ contains an element $\omega$ of multiplicative order $9$, and consider the following irreducible $3$-dimensional representations of $G$:
@@
One can check by basic linear algebra that $\psi_{j}(g)$ has the eigenvalue $1$ for some $g\in G$ and $j\in\{1,2\}$ if and only if $g$ belongs to one of the order $3$ subgroups $\langle b\rangle$, $\langle a^{3}b\rangle$, $\langle a^{6}b\rangle$ of $G$. Computing the corresponding eigenvectors we obtain the result. ∎
@@
So $x_{1}(w_{1})=x_{2}(w_{1})$ or $x_{1}(w_{1})=\varepsilon x_{2}(w_{1})$ or $x_{1}(w_{1})=\varepsilon^{2}x_{2}(w_{1})$. Taking into account (33) and (35) we conclude that $w_{1}$ is a non-zero scalar multiple of $[1,1,1]^{T}$,$[\varepsilon^{2},\varepsilon,1]^{T}$, or $[\varepsilon,\varepsilon^{2},1]^{T}\}$. By Lemma 8.12 it follows that $\mathrm{Stab}_{G}(w_{1})\in\{\langle b\rangle,\langle a^{3}b\rangle,\langle a^{6}b\rangle\}$. In particular, we are done if $w_{2}=0$, since then $\mathrm{Stab}_{G}(w_{1},w_{2})=\mathrm{Stab}_{G}(w_{1})$.
@@
Now (38), (39) clearly imply that $w_{2}$ is a non-zero scalar multiple of $[1,1,1]^{T}$, $[\varepsilon,\varepsilon^{2},1]^{T}$, or $[\varepsilon^{2},\varepsilon,1]^{T}$, implying in turn by Lemma 8.12 that $\mathrm{Stab}_{G}(w_{2})\in\{\langle b\rangle,\langle a^{3}b\rangle,\langle a^{6}b\rangle\}$.
@@
Thus we showed that both $w_{1}$ and $w_{2}$ are scalar multiples of $[1,1,1]^{T}$, $[\varepsilon^{2},\varepsilon,1]^{T}$, or $[\varepsilon,\varepsilon^{2},1]^{T}$. If $w_{1}=0$ or $w_{2}=0$, then we have the desired $\mathrm{Stab}_{G}(w_{1},w_{2})\neq\{1_{G}\}$ by Lemma 8.12. If both $w_{1}$ and $w_{2}$ are non-zero scalar multiples of $[1,1,1]^{T}$, $[\varepsilon^{2},\varepsilon,1]^{T}$, or $[\varepsilon,\varepsilon^{2},1]^{T}$, then again using Lemma 8.12 one can verify by direct computation that for such a pair $(w_{1},w_{2})$, we have
@@
$0=f_{(1,\varepsilon)}^{(3)}(w_{1})=x_{1}x_{2}^{2}+\varepsilon^{2}x_{2}(x_{1}x_{2})+\varepsilon x_{3}x_{1}^{2})(w_{1})=\varepsilon x_{1}(w_{1})(x_{1}x_{3}-x_{2}^{2})(w_{1}).$
Thus $(x_{1}x_{3})(w_{1})=x_{2}^{2}(w_{1})$, and this together with $(x_{1}x_{2})(w_{1})=x_{3}^{2}(w_{1})$ implies that $w_{1}$ is a non-zero scalar multiple of $[1,1,1]^{T}$, $[\varepsilon,\varepsilon^{2},1]^{T}$, or $[\varepsilon^{2},\varepsilon,1]^{T}$.
If $x_{3}^{3}(w_{1})=x_{1}^{3}(w_{1})$ (the other alternative from (41), then (recall that $x_{2}^{3}(w_{1})=x_{1}^{3}(w_{1})$) there exist some cubic roots $\nu_{2},\nu_{3}$ of $1$ such that $w_{1}$ is a non-zero scalar multiple of $[1,\nu_{2},\nu_{3}]$. Then $0=f_{(1,\varepsilon)}^{(3)}(w_{1})$ reduces to
@@
The sum of three cubic roots of $1$ is zero only if the three summands are $1,\varepsilon,\varepsilon^{2}$ in some order. We conclude that $w_{1}$ is a non-zero scalar multiple of $[1,1,1]^{T}$, $[\varepsilon,\varepsilon^{2},1]^{T}$, or $[\varepsilon^{2},\varepsilon,1]^{T}$.
@@
and so $y_{1}^{3}(w_{2})=y_{2}^{3}(w_{2})$. Obviously, this together with $y_{2}^{2}(w_{2})=y_{1}(w_{2})y_{3}(w_{2})$ means that $w_{2}$ is a non-zero scalar multiple of $[1,1,1]^{T}$, $[\varepsilon,\varepsilon^{2},1]^{T}$, or $[\varepsilon^{2},\varepsilon,1]^{T}$.
@@
The map $a\mapsto a^{2}$, $b\mapsto b$ extends to an automorphism $\alpha$ of $G$, and for $\chi=(\varepsilon,\varepsilon)$ we have $\chi\circ\alpha=(\varepsilon^{2},\varepsilon)$, whereas for $\chi=(\varepsilon,1)$ we have $\chi\circ\alpha=(\varepsilon^{2},1)$. So by Lemma 3.11, statement (ii) holds also for the weight $\chi=(\varepsilon^{2},\varepsilon)$ and $\chi=(\varepsilon^{2},1)$. The only information on $\varepsilon$ used in the constructions of relative invariants above was that it has multiplicative order $3$. Therefore we can replace it by the other element of multiplicative order $3$, namely by $\varepsilon^{2}$, and we get that (ii) holds also for the weights $\chi\in\{\varepsilon,\varepsilon^{2}),(\varepsilon^{2},\varepsilon^{2}),(1,\varepsilon^{2})\}$. ∎
@@
(42) $f(v)=f(v^{\prime})\text{ for all homogeneous }f\in K[W_{1}\oplus W_{2}\oplus U]^{G}\text{ where }\deg(f)\leq 10.$
We need to show that $G\cdot v=G\cdot v^{\prime}$. Replacing $v^{\prime}$ by an appropriate element in its $G$-orbit, we may assume by Proposition 8.11 that $(w_{1},w_{2})=(w^{\prime}_{1},w^{\prime}_{2})$. Moreover, $G\cdot u=G\cdot u^{\prime}$, since $\mathsf{D}(G/G^{\prime})=\mathsf{D}(\mathrm{C}_{3}\times\mathrm{C_{3}})=5$. Therefore it is sufficient to deal with the case when $(w_{1},w_{2})\in W_{1}\oplus W_{2}$ is non-zero.
@@
Case II.: $\mathrm{Stab}_{G}(w_{1},w_{2})=\{1_{G}\}$. We claim that $u=u^{\prime}$; that is, we claim that $u_{\chi}=u^{\prime}_{\chi}$ for all $\chi\in\widehat{G}$. This is obvious for $\chi=(1,1)\in\widehat{G}$, since $u$ and $u^{\prime}$ have the same $G$-orbit. For $\chi\in\widehat{G}\setminus\{(1,1)\}$, by Lemma 8.13 there exists an $f\in K[W_{1}\oplus W_{2}\oplus U]^{G,\chi^{-1}}$ with $\deg(f)\leq 9$ and $f(w_{1},w_{2})\neq 0$. Then $ft_{\chi}$ is a $G$-invariant of degree at most $10$. Thus by (42) we have $(ft_{\chi})(w_{1},w_{2},u)=(ft_{\chi})(w_{1},w_{2},u^{\prime})$, implying in turn that $t_{\chi}(u)=t_{\chi}(u^{\prime})$, i.e. $u_{\chi}=u^{\prime}_{\chi}$. This finishes the proof of the inequality $\beta_{\mathrm{sep}}^{K}(G)\leq 10$.
To see the reverse inequality, consider the $G$-module $W_{1}\oplus U_{(1,\varepsilon^{2})}$. Consider the vectors $v:=([1,0,0]^{T},\varepsilon)$ and $v^{\prime}:=([1,0,0]^{T},\varepsilon^{2})$. We have $(f_{(1,\varepsilon)}^{(4)}t_{(1,\varepsilon^{2})})(v)=\varepsilon$, whereas $(f_{(1,\varepsilon)}^{(4)}t_{(1,\varepsilon^{2})})(v^{\prime})=\varepsilon^{2}$, so the invariant $(f_{(1,\varepsilon)}^{(4)}t_{(1,\varepsilon^{2})}$) separates $v$ and $v^{\prime}$. We claim that no $G$-invariant of degree at most $9$ separates $v$ and $v^{\prime}$. The elements of $K[W_{1}]$ and $K[U_{(1,\varepsilon^{2})}]^{G}=K[t_{(1,\varepsilon)}^{3}]$ agree on $v$ and $v^{\prime}$. Suppose for contradiction that there exists a multihomogeneous invariant $f=ht_{(1,\varepsilon^{2})}$ (respectively $f=ht_{(1,\varepsilon^{2})}^{2}$) of degree at most $9$ with $f(v)\neq f(v^{\prime})$, where $h\in K[W_{1}]^{G,(1,\varepsilon)}$ (respectively $h\in K[W_{1}]^{G,(1,\varepsilon^{2})}$). Then $h$ has a monomial of the form $x^{d}$ with non-zero coefficients (since $0=x_{2}(v)=x_{3}(v)=x_{2}(v^{\prime})=x_{3}(v^{\prime})$). Moreover, $x_{1}^{d}$ must be an $\langle a\rangle$-invariant monomial. However, the smallest $d$ for which $x_{1}^{d}$ is $\langle a\rangle$-invariant is $9$. This is a contradiction, because the degree of $h$ is strictly less than $9$. The inequality $\beta_{\mathrm{sep}}(G,W_{1}\oplus U_{(1,\varepsilon)})\geq 10$ is proved. ∎
Acknowledgements
This research was partially supported by the Hungarian National Research, Development and Innovation Office, NKFIH K 138828.

### v3

@@ 7. The group $\mathrm{M}_{27}$
is the non-abelian group of order $27$ with an index $3$ cyclic subgroup. Assume that $K$ contains an element $\omega$ of multiplicative order $9$, and consider the following two non-isomorphic irreducible $3$-dimensional representations of $G$:
@@
One can check by basic linear algebra that $\psi_{j}(g)$ has the eigenvalue $1$ for some $g\in G$ and $j\in\{1,2\}$ if and only if $g$ belongs to one of the order $3$ subgroups $\langle b\rangle$, $\langle a^{3}b\rangle$, $\langle a^{6}b\rangle$ of $G$. By computing the corresponding eigenvectors we obtain the result. ∎
@@
So $x_{1}(w_{1})=x_{2}(w_{1})$ or $x_{1}(w_{1})=\varepsilon x_{2}(w_{1})$ or $x_{1}(w_{1})=\varepsilon^{2}x_{2}(w_{1})$. Taking into account (23) and (25) we conclude that $w_{1}$ is a non-zero scalar multiple of $1,1,1]^{T}$$\varepsilon^{2},\varepsilon,1]^{T}$, or $\varepsilon,\varepsilon^{2},1]^{T}\}$. By Lemma 7.2 it follows that $\mathrm{Stab}_{G}(w_{1})\in\{\langle b\rangle,\langle a^{3}b\rangle,\langle a^{6}b\rangle\}$. In particular, we are done if $w_{2}=0$, since then $\mathrm{Stab}_{G}(w_{1},w_{2})=\mathrm{Stab}_{G}(w_{1})$.
@@
Now (28), (29) clearly imply that $w_{2}$ is a non-zero scalar multiple of $1,1,1]^{T}$, $\varepsilon,\varepsilon^{2},1]^{T}$, or $\varepsilon^{2},\varepsilon,1]^{T}$, implying in turn by Lemma 7.2 that $\mathrm{Stab}_{G}(w_{2})\in\{\langle b\rangle,\langle a^{3}b\rangle,\langle a^{6}b\rangle\}$.
@@
Thus we showed that both $w_{1}$ and $w_{2}$ are scalar multiples of $1,1,1]^{T}$, $\varepsilon^{2},\varepsilon,1]^{T}$, or $\varepsilon,\varepsilon^{2},1]^{T}$. If $w_{1}=0$ or $w_{2}=0$, then we have the desired $\mathrm{Stab}_{G}(w_{1},w_{2})\neq\{1_{G}\}$ by Lemma 7.2. If both $w_{1}$ and $w_{2}$ are non-zero scalar multiples of $1,1,1]^{T}$, $\varepsilon^{2},\varepsilon,1]^{T}$, or $\varepsilon,\varepsilon^{2},1]^{T}$, then again using Lemma 7.2 one can verify by direct computation that for such a pair $(w_{1},w_{2})$, we have
@@
$0=f_{(1,\varepsilon)}^{(3)}(w_{1})=(x_{1}x_{2}^{2}+\varepsilon^{2}x_{2}(x_{1}x_{2})+\varepsilon x_{3}x_{1}^{2})(w_{1})=\varepsilon x_{1}(w_{1})(x_{1}x_{3}-x_{2}^{2})(w_{1}).$
Thus $(x_{1}x_{3})(w_{1})=x_{2}^{2}(w_{1})$, and this together with $(x_{1}x_{2})(w_{1})=x_{3}^{2}(w_{1})$ implies that $w_{1}$ is a non-zero scalar multiple of $1,1,1]^{T}$, $\varepsilon,\varepsilon^{2},1]^{T}$, or $\varepsilon^{2},\varepsilon,1]^{T}$.
If $x_{3}^{3}(w_{1})=x_{1}^{3}(w_{1})$ (the other alternative from (31), then (recall that $x_{2}^{3}(w_{1})=x_{1}^{3}(w_{1})$) there exist some cubic roots $\nu_{2},\nu_{3}$ of $1$ such that $w_{1}$ is a non-zero scalar multiple of $1,\nu_{2},\nu_{3}]$. Then $0=f_{(1,\varepsilon)}^{(3)}(w_{1})$ reduces to
@@
The sum of three cubic roots of $1$ is zero only if the three summands are $1,\varepsilon,\varepsilon^{2}$ in some order. We conclude that $w_{1}$ is a non-zero scalar multiple of $1,1,1]^{T}$, $\varepsilon,\varepsilon^{2},1]^{T}$, or $\varepsilon^{2},\varepsilon,1]^{T}$.
@@
and so $y_{1}^{3}(w_{2})=y_{2}^{3}(w_{2})$. Obviously, this together with $y_{2}^{2}(w_{2})=y_{1}(w_{2})y_{3}(w_{2})$ means that $w_{2}$ is a non-zero scalar multiple of $1,1,1]^{T}$, $\varepsilon,\varepsilon^{2},1]^{T}$, or $\varepsilon^{2},\varepsilon,1]^{T}$.
@@
Therefore $\mathrm{Stab}_{G}(w_{1},w_{2})=\mathrm{Stab}_{G}(w_{1})\neq\{1_{G}\}$. This finishes the proof for Case II.
The map $a\mapsto a^{2}$, $b\mapsto b$ extends to an automorphism $\alpha$ of $G$, and for $\chi=(\varepsilon,\varepsilon)$ we have $\chi\circ\alpha=(\varepsilon^{2},\varepsilon)$. So by Lemma 2.5 (iii), the conclusion of statement (ii) holds also for the weight $\chi=(\varepsilon^{2},\varepsilon)$. The only information on $\varepsilon$ used in the constructions of relative invariants in Cases I and II was that it has multiplicative order $3$. Therefore we can replace it by the other element of multiplicative order $3$, namely by $\varepsilon^{2}$, and we get that (ii) holds also for the weights $\chi\in\{\varepsilon,\varepsilon^{2}),(\varepsilon^{2},\varepsilon^{2}),(1,\varepsilon^{2})\}$. ∎
@@
(32) f(v)=f(v) $f(v^{\prime})\text{ for all homogeneous }f\in K[W_{1}\oplus W_{2}\oplus U]^{G}\text{ where }\deg(f)\leq 10$.
We need to show that $G\cdot v=G\cdot v^{\prime}$. Replacing $v^{\prime}$ by an appropriate element in its $G$-orbit, we may assume by Proposition 7.1 that $(w_{1},w_{2})=(w^{\prime}_{1},w^{\prime}_{2})$. Moreover, $G\cdot u=G\cdot u^{\prime}$, since $\mathsf{D}(G/G^{\prime})=\mathsf{D}(\mathrm{C}_{3}\times\mathrm{\mathrm{C}_{3}})=5$. Therefore it is sufficient to deal with the case when $(w_{1},w_{2})\in W_{1}\oplus W_{2}$ is non-zero.
@@
[…]
To see the reverse inequality, consider the $KG$-module $W_{1}\oplus U_{(1,\varepsilon^{2})}$. Consider the vectors $v:=([1,0,0]^{T},\varepsilon)$ and $v^{\prime}:=([1,0,0]^{T},\varepsilon^{2})$. We have $(f_{(1,\varepsilon)}^{(4)}t_{(1,\varepsilon^{2})})(v)=\varepsilon$, whereas $(f_{(1,\varepsilon)}^{(4)}t_{(1,\varepsilon^{2})})(v^{\prime})=\varepsilon^{2}$, so the invariant $(f_{(1,\varepsilon)}^{(4)}t_{(1,\varepsilon^{2})}$) separates $v$ and $v^{\prime}$. We claim that no $G$-invariant of degree at most $9$ separates $v$ and $v^{\prime}$. The elements of $K[W_{1}]$ and $K[U_{(1,\varepsilon^{2})}]^{G}=K[t_{(1,\varepsilon)}^{3}]$ agree on $v$ and $v^{\prime}$. Suppose for contradiction that there exists a multihomogeneous invariant $f=ht_{(1,\varepsilon^{2})}$ (respectively $f=ht_{(1,\varepsilon^{2})}^{2}$) of degree at most $9$ with $f(v)\neq f(v^{\prime})$, where $h\in K[W_{1}]^{G,(1,\varepsilon)}$ (respectively $h\in K[W_{1}]^{G,(1,\varepsilon^{2})}$). Then $h$ has a monomial of the form $x_{1}^{d}$ with non-zero coefficients (since $0=x_{2}(v)=x_{3}(v)=x_{2}(v^{\prime})=x_{3}(v^{\prime})$). Moreover, $x_{1}^{d}$ must be an $\langle a\rangle$-invariant monomial. However, the smallest $d$ for which $x_{1}^{d}$ is $\langle a\rangle$-invariant is $9$. This is a contradiction, because the degree of $h$ is strictly less than $9$. The inequality $\beta_{\mathrm{sep}}(G,W_{1}\oplus U_{(1,\varepsilon)})\geq 10$ is proved. ∎